OpenStudy (anonymous):

A pizza shop offers the toppings shown below. How many different 3-topping pizzas can you make?pepperoni mushrooms sausage onion ham (1 point) a) 6 b)10 c)4 d)5

OpenStudy (anonymous):

how many ways can you choose 3 items from a set of 5? i.e. what is \(_5C_3\) ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

5 ?

geerky42 (geerky42):

\[\Large ^nC_r = \dfrac{n!}{(n-r)!r!}\] So \(\Large^5C_3 = \dfrac{5!}{(5-3)!3!}\)

OpenStudy (anonymous):

i am lost :(

geerky42 (geerky42):

5! = 5·4·3·2·1 3! = 3·2·1 2! = 2·1 So \(\dfrac{5!}{3!2!} = \dfrac{5\cdot4\cdot3\cdot2\cdot1}{(3\cdot2\cdot1)\cdot(2\cdot1)} = \dfrac{5\cdot4}{2\cdot1} = \cdots~?\)